OpenStudy (anonymous):

What is the exact value of the expression the square root 72. − the square root of 8. + the square root of 128.? Simplify if possible.

hartnn (hartnn):

\(\sqrt{72}-\sqrt{8}+\sqrt{128}\) this , right ? can you factor, 72 =.... ? 8=...? 128 =...?

OpenStudy (anonymous):

72 = 6 sqrt of 2 8 = 2 sqrt of 2 128 = 8 sqrt 2

OpenStudy (tkhunny):

Well, add them up.

hartnn (hartnn):

good! now just factor out sqrt 2 from each term :) can u ?

OpenStudy (anonymous):

So It would look like ( 6-2+8) sqrt of 2 ?

hartnn (hartnn):

exactly!

OpenStudy (anonymous):

which would be 12sqrt of 2

hartnn (hartnn):

that would be correct :) you are good at this! :)

OpenStudy (anonymous):

Thank you but dividing is my problem actually .

hartnn (hartnn):

any particular problem ?

OpenStudy (anonymous):